# Homological shifts of a complementary edge ideal

### v1

@@ 3. Projective dimensions functions
    $i\in\mathrm{set}(\beta)$ if and only if $j\neq n$. In particular, $\mathrm{set}(\beta)\subseteq[n-2]$.
@@
    If $j\in[n-2]$, then $j\in\mathrm{set}(\beta)$ if and only if there is some $2\leq t\leq s$ such that $u_{t}=x_{j}x_{k}$ and $k>j$.
@@
(1) If $j\neq n$, then there is some $k\in[n]$ such that $k>j$ and $k$ is adjacent to $j$. Since $j\geq i$, we have $k>i$. By Lemma 3.3, $i\in\mathrm{set}(\beta)$.
If $j=n$, then i=n1i=n-1, since $n$ is a leaf of $G$. If $(n-1)\in\mathrm{set}(\beta)$, then by Lemma 3.3, $n$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$. Thus there exists a closed walk $y_{1}-y_{2}-\cdots-y_{2\ell+1}-y_{2\ell+2}$ such that $y_{1}=y_{2\ell+2}=n$. However, since $G$ is a tree, every edge of $G$ appears an even number of times in this walk. Therefore, the walk is of even length. However, this walk has length $2\ell+1$, which is a contradiction. This implies $i\notin\mathrm{set}(\beta)$.
(2) If $j\in\mathrm{set}(\beta)$, then, by Lemma 3.3, there is some $k^{\prime}\in[n]$ such that $k^{\prime}>j$ and either $\{i,k^{\prime}\}\in E(G)$ or $i$ is even-connected to $k^{\prime}$ with respect to (s1)(s-1)-fold product $u_{2}\cdots u_{s}$. Since $G$ is a tree, based on the labeling of $G$’s vertices as described in Remark 3.5, we know that $j$ is the unique vertex satisfying $\{j,i\}\in E(G)$ and $j>i$. Thus $i$ is even-connected to $k^{\prime}$ with respect to (s1)(s-1)-fold product $u_{2}\cdots u_{s}$ by the walk $y_{1}-y_{2}-\cdots-y_{2\ell+1}-y_{2\ell+2}$. Let $y_{1}=i$ and $y_{2\ell+2}=k^{\prime}$, and let $p$ be the largest integer satisfying $y_{p}=i$. If $p\neq 1$, then $y_{1}-y_{2}-\cdots-y_{p-1}-y_{p}$ is a closed walk. This implies that $p$ is odd and $x_{y_{p+1}}x_{y_{p+2}}$ belongs to $\{u_{2},\ldots,u_{s}\}$. Again based on the labeling of $G$’s vertices as described in Remark 3.5, we know that $j=y_{p+1}$. If $y_{p+2}>j$, we can set $k=y_{p+2}$ and we are done. If $y_{p+2}<j$, let $q$ be the largest integer satisfying $y_{q}=j$ and $q\geq p+1$. Then q(p+1)q-(p+1) is even and $y_{q+1}>j$, since $G$ is a tree. We put $k=y_{p+1}$ and the result follows. The converse conclusion follows from (1).    ∎
Let $G$ be a graph and $e=\{i,j\}$ be an edge of $G$, we set $\min(e)=\min\{i,j\}$.
@@
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$, where $e_{n-1}=\{n-1,n\}=x_{n-1}x_{n}$. Then $\min(e_{1}),\ldots,\min(e_{n-2})$ are pairwise distinct.
For any $s\in[n-2]$, we choose $\beta_{s}=\frac{\alpha^{s}}{e_{1}\cdots e_{s}}$. Then $\beta_{s}\in\mathcal{G}(I_{c}(G)^{s})$ and, by Lemma 3.6.(1), we can deduce that
$\{\min(e_{1}),\ldots,\min(e_{s})\}\subseteq\mathrm{set}(\beta_{s}).$
This implies that $\mathrm{pd}(I_{c}(G)^{s})\geq s$.
For any $\beta\in\mathcal{G}(I_{c}(G)^{s})$, we can write $\beta$ as $\beta=\frac{\alpha^{s}}{e_{i_{1}}\cdots e_{i_{s}}}$, where each $i_{j}\in[n-1]$. If $i\in\mathrm{set}(\beta)$, then by Lemma 3.6 (2), there is a $t\in[s]$ such that $i=\min(u_{t})$. This implies
(3) $\mathrm{set}(\beta)=\{\min(e_{i_{j}}))\mid j\in[s]\}\setminus\{n-1\}.$
Therefore, $|\mathrm{set}(\beta)|\leq s$. It follows from the arbitrariness of $\beta$ that $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\operatorname{set}(\beta)|:\ \beta\in\mathcal{G}(I_{c}(G)^{s})\}\leq s$.
Suppose that $s\geq n-2$. Then
$n-2=\mathrm{pd}(I_{c}(G)^{n-2})\leq\mathrm{pd}(I_{c}(G)^{s})\leq n-2.$
This implies $\mathrm{pd}(I_{c}(G)^{s})=n-2$, as desired.    ∎
This result generalizes [17, Proposition 4.6], which proves that the same formula holds for paths, a specific type of trees.
The projective dimension of $I_{c}(G)$ for all connected graph is given.
@@
If $G$ is a tree, then $\mathrm{pd}(I_{c}(G))=1$ by Theorem 3.7. If $G$ is not a tree, there exists a cycle $C$ that is a subgraph of $G$. By Lemma 3.4 and Lemma 1.3, we have $\mathrm{pd}(I_{c}(G))\geq\mathrm{pd}(I_{c}(C))=2$.    ∎
For any connected graph $G$, it is known that $\mathrm{pd}(I_{c}(G)^{s})$ is non-decreasing and $\mathrm{depth}(I_{c}(G)^{s})$ is non-increasing in $s$. This can be improved upon when $G$ is bipartite.
@@
Let $G$ be a connected bipartite graph with $n$ vertices. Then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches its maximum value n2n-2.
@@
By [17, Proposition 4.3] and the Auslander-Buchsbaum formula, we know that $\mathrm{pd}(I_{c}(G)^{s})=n-2$ for all $s\gg 0$. Suppose that $\mathrm{pd}(I_{c}(G)^{s})=k\leq n-3$. We can choose $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ such that $|\operatorname{set}(\beta)|=k$. Let $\ell$ be the minimal element of $n]\setminus\operatorname{set}(\beta)$, and let $T$ be a spanning tree of $G$, and label its vertices as described in Remark 3.5. Since $k\leq n-3$, it follows that $\ell\leq n-2$. Thus, there exists an integer $m$ with $\ell<m\leq n-1$ such that $\{\ell,m\}\in E(T)$. Set $u_{s+1}=x_{\ell}x_{m}$ and $\gamma=\beta\cdot\frac{\alpha}{u_{s+1}}$. By Lemma 3.3, $\ell\in\operatorname{set}(\gamma)$. Therefore, $|\operatorname{set}(\gamma)|\geq k+1$. It follows that $\mathrm{pd}(I_{c}(G)^{s+1})\geq k+1$, which completes the proof.    ∎
If $G$ is a connected non-bipartite graph with $n$ vertices, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ is n1n-1. Using the same method as in Lemma 3.9, we can prove that $\mathrm{pd}(I_{c}(G)^{s})$ is strictly increasing in $s$ until it reaches n2n-2. However, it is difficult to prove that it continues to increase strictly when it reaches n2n-2.
By [4], we know that the limit $\lim\limits_{k\rightarrow\infty}\mathrm{depth}(S/I^{k})$ exists for any ideal $I\subset S$, that is, $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k+1})$ for all $k\gg 0$. The least integer $k_{0}>0$ for which $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k_{0}})$ for all $k\geq k_{0}$ is called the index of depth stability of $I$ and is denoted by $\mathrm{dstab}(I)$.
@@
By Theorem 3.7 together with the Auslander-Buchsbaum formula, we can show that if $G$ is a tree, then $\mathrm{dstab}(I_{c}(G))=n-2$.
@@
Suppose that $G$ contains at least one cycle, denoted $C$. It suffices to show that $\mathrm{dstab}(I_{c}(G))<n-2$. By Lemma 3.4, $\mathrm{pd}(I_{c}(C))=2$. By Lemma 1.3, it follows that $\mathrm{pd}(I_{c}(G))\geq\mathrm{pd}(I_{c}(C))=2$. Assume, on the contrary, that $\mathrm{dstab}(I_{c}(G))\geq n-2$. Then, according to Proposition 3.9, we have
$\mathrm{pd}(I_{c}(G)^{n-2})\geq\mathrm{pd}(I_{c}(G)^{n-3})+1\geq\cdots\geq\mathrm{pd}(I_{c}(G))+(n-3)\geq 2+(n-3)=n-1.$
This yields a contradiction, completing the proof.    ∎
4. Homological shift algebras
In this section, we will examine the following conjecture and prove that it is true for graphs that are either trees or cycles.
Conjecture 4.1.
Let $G$ be a connected graph, and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i$.

### v5

@@ 4. Projective dimension functions
    $i\in\mathrm{set}(\beta)$ if and only if $j\neq n$.
@@
    If $j\in\mathrm{set}(\beta)$, then there exists some $2\leq t\leq s$ such that $u_{t}=x_{j}x_{k}$ and $k>j$.
  3. (3)
    $\mathrm{set}(\beta)\subseteq[n-2]$. In particular, $\mathrm{pd}(I_{c}(G)^{s})\leq n-2$ for all $s\geq 1$.
@@
(1) If $j\neq n$, then there exists $k\in[n]$ such that $k>j$ and $k$ is adjacent to $j$. Since $j>i$, we have $k>i$; by Lemma 4.3, $i\in\mathrm{set}(\beta)$.
If $j=n$, then i=n1i=n-1. Suppose for contradiction that $n-1\in\mathrm{set}(\beta)$. By Lemma 4.3, $n$ is even-connected to itself with respect to the product $u_{2}\cdots u_{s}$, so there exists a closed walk $y_{1}-y_{2}-\cdots-y_{2\ell+2}$ with $y_{1}=y_{2\ell+2}=n$. Since $G$ is a tree (hence acyclic), every edge in this closed walk must appear an even number of times, which implies the walk has even length. But the walk has length $2\ell+1$ (odd), a contradiction. Thus $i=n-1\notin\mathrm{set}(\beta)$.
(2) If $j\in\mathrm{set}(\beta)$, then by Lemma 4.3, there exists $k^{\prime}\in[n]$ with $k^{\prime}>j$ such that either $\{i,k^{\prime}\}\in E(G)$ or $i$ is even-connected to $k^{\prime}$ with respect to $u_{2}\cdots u_{s}$. By the vertex labeling in Remark 4.5, $j$ is the unique neighbor of $i$ satisfying $j>i$. Given $k^{\prime}>j$, we have $\{i,k^{\prime}\}\notin E(G)$. Hence, $i$ must be even-connected to $k^{\prime}$ via a walk $y_{1}-y_{2}-\cdots-y_{2\ell+2}$ with $y_{1}=i$ and $y_{2\ell+2}=k^{\prime}$. Let $p$ be the largest integer such that $y_{p}=i$.
If $p\neq 1$, then $y_{1}-y_{2}-\cdots-y_{p}$ is a closed walk, so $p$ is odd. It follows that $x_{y_{p+1}}x_{y_{p+2}}\in\{u_{2},\ldots,u_{s}\}$, and Lemma 4.6 gives $y_{p+1}=j$.
- If $y_{p+2}>j$, then $x_{y_{p+1}}x_{y_{p+2}}=x_{j}x_{y_{p+2}}$ is the desired edge.
- If $y_{p+2}<j$, let $q$ denote the maximal integer such that $y_{q}=j$ and $q\geq p+1$. As all closed walks in $G$ have even length, q(p+1)q-(p+1) is even; moreover, $y_{q+1}>j$ by Lemma 4.6. The fact that q(p+1)q-(p+1) is even implies $x_{y_{q}}x_{y_{q+1}}\in\{u_{2},\dots,u_{s}\}$. Thus, $x_{y_{q}}x_{y_{q+1}}=x_{j}x_{y_{q+1}}$ is the required edge.
(3) We first note $n\notin\mathrm{set}(\beta)$ by Lemma 4.3. It remains to show $n-1\notin\mathrm{set}(\beta)$. Assume for contradiction that $n-1\in\mathrm{set}(\beta)$. By part (2), there exists $t\in[s]$ such that $u_{t}=x_{n-1}x_{n}$. But by part (1), since $u_{t}$ has $j=n$, we get $n-1\notin\mathrm{set}(\beta)$, a contradiction. Hence $\mathrm{set}(\beta)\subseteq[n-2]$.
Let $G$ be a graph on $[n]$, and $e=\{i,j\}$ be an edge of $G$, we set $\min(e)=\min\{i,j\}$.
@@
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$, where $e_{n-1}=\{n-1,n\}=x_{n-1}x_{n}$. Then, $\min(e_{1}),\min(e_{2}),\ldots,\min(e_{n-2})$ are pairwise distinct. In fact, if $\min(e_{i})=\min(e_{j}):=k$, then $e_{i}=e_{j}=\{k,\ell\}$, where $\ell$ is the unique parent of $k$, see Remark 4.5.
First, consider $s\in[n-2]$. Choose $\beta_{s}=\frac{\alpha^{s}}{e_{1}e_{2}\cdots e_{s}}$. It is clear that $\beta_{s}\in\mathcal{G}(I_{c}(G)^{s})$. By Lemma 4.7 (1), we directly deduce that
$\{\min(e_{1}),\min(e_{2}),\ldots,\min(e_{s})\}\subseteq\operatorname{set}(\beta_{s}),$
which implies $\operatorname{pd}(I_{c}(G)^{s})\geq s$ by formula (2).
Next, take an arbitrary $\beta\in\mathcal{G}(I_{c}(G)^{s})$. Then $\beta$ can be written as $\beta=\frac{\alpha^{s}}{e_{i_{1}}e_{i_{2}}\cdots e_{i_{s}}}$, where $i_{j}\in[n-1]$ for each $j\in[s]$. If $i\in\operatorname{set}(\beta)$, then by Lemma 4.7 (2), there exists some $t\in[s]$ such that $i=\min(e_{i_{t}})$. Combining this with Lemma 4.7 (1) and (3), we immediately deduce that
(3) $\operatorname{set}(\beta)=\{\min(e_{i_{j}})\mid j\in[s]\}\setminus\{n-1\}.$
Therefore, $|\operatorname{set}(\beta)|\leq s$. Since $\beta$ is arbitrary, substituting into formula (2) gives
$\operatorname{pd}(I_{c}(G)^{s})=\max\left\{|\operatorname{set}(\beta)|:\beta\in\mathcal{G}(I_{c}(G)^{s})\right\}\leq s.$
Combining the previously established lower bound ($\operatorname{pd}(I_{c}(G)^{s})\geq s$) and the upper bound above, we conclude $\operatorname{pd}(I_{c}(G)^{s})=s$ for $s\in[n-2]$.
Now, consider $s\geq n-2$. By Lemma 2.6, we have $\operatorname{pd}(I_{c}(G)^{n-2})\leq\operatorname{pd}(I_{c}(G)^{s})$. On the other hand, we have already proven that $\operatorname{pd}(I_{c}(G)^{n-2})=n-2$, and Lemma 4.7 (3) further implies $\operatorname{pd}(I_{c}(G)^{s})\leq n-2$ for all $s\geq 1$. Thus,
$n-2=\operatorname{pd}(I_{c}(G)^{n-2})\leq\operatorname{pd}(I_{c}(G)^{s})\leq n-2,$
which forces $\operatorname{pd}(I_{c}(G)^{s})=n-2$. This completes the proof.
This result generalizes [14, Proposition 4.6], which proves that the same formula holds for paths, a specific type of trees.
We next determine the projective dimension of $I_{c}(G)$ for all connected graphs $G$.
@@
If $G$ is a tree, then $\mathrm{pd}(I_{c}(G))=1$ by Theorem 4.8. If $G$ is not a tree, there exists a cycle $C$ that is a subgraph of $G$. By Theorem 4.4 and Theorem 2.5, we have $\mathrm{pd}(I_{c}(G))\geq\mathrm{pd}(I_{c}(C))=2$.
For any connected graph $G$, it is known from Lemma 2.5 that $\mathrm{pd}(I_{c}(G)^{s})$ is a non-decreasing function of $s$. This result can be strengthened. To do so, we first recall the definitions of a spanning subgraph and a bipartite graph: a subgraph $H$ of $G$ is spanning if they share the same vertex set, i.e., $V(H)=V(G)$; a bipartite graph is a graph whose vertex set can be partitioned into two disjoint non-empty subsets $X$ and $Y$ such that every edge connects a vertex in $X$ to a vertex in $Y$.
@@
Let $G$ be a connected graph with $n$ vertices. Then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches the value n2n-2.
In particular, if we furhter assume that $G$ is bipartite, then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches the maximal value.
@@
Let $T$ be a spanning tree of $G$, and label its vertices as described in Remark 4.5. Note that this also defines a labeling on $V(G)$ that satisfies the FM condition. Suppose that $\mathrm{pd}(I_{c}(G)^{s})=k\leq n-3$. We can choose $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ such that $|\operatorname{set}(\beta)|=k$. Let $\ell$ be the minimal element of $n]\setminus\operatorname{set}(\beta)$. Since $k\leq n-3$, it follows that $\ell\leq n-2$. Thus, there exists an integer $m$ with $\ell<m\leq n-1$ such that $\{\ell,m\}\in E(T)$. Set $u_{s+1}=x_{\ell}x_{m}$ and $\gamma=\beta\cdot\frac{\alpha}{u_{s+1}}$. By Lemma 4.3, $\ell\in\operatorname{set}(\gamma)$. Therefore, $|\operatorname{set}(\gamma)|\geq k+1$. It follows that $\mathrm{pd}(I_{c}(G)^{s+1})\geq k+1$, which completes the proof of the first assertion.
For the second assertion, we only need to note that if $G$ is a connected bipartite graph, then $\max\{\mathrm{pd}(I_{c}(G)^{s})|s\geq 1\}=n-2$, by [14, Proposition 4.3] and the Auslander-Buchsbaum formula.
If $G$ is a connected non-bipartite graph with $n$ vertices, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ is n1n-1. However, we have not yet established whether it continues to increase strictly upon reaching the value n2n-2.
By [4], we know that $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k+1})$ for all $k\gg 0$. The least integer $k_{0}>0$ for which $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k_{0}})$ for all $k\geq k_{0}$ is called the index of depth stability of $I$ and is denoted by $\mathrm{dstab}(I)$.
@@
Combining Theorem 4.8 with the Auslander-Buchsbaum formula, we conclude that if $G$ is a tree, then $\mathrm{dstab}(I_{c}(G))=n-2$.
@@
Suppose $G$ contains at least one cycle. It suffices to show that $\mathrm{dstab}(I_{c}(G))<n-2$. By Proposition 4.9, we have $\mathrm{pd}(I_{c}(G))=2$. Assume for contradiction that $\mathrm{dstab}(I_{c}(G))\geq n-2$. Then, by Proposition 4.10, the following chain of inequalities holds:
$\mathrm{pd}(I_{c}(G)^{n-2})\geq\mathrm{pd}(I_{c}(G)^{n-3})+1\geq\cdots\geq\mathrm{pd}(I_{c}(G))+(n-3)=n-1.$
This is a contradiction, which completes the proof.
We now consider connected non-bipartite graphs. First, recall that a unicyclic graph is a connected graph containing precisely one cycle. Note that every connected non-bipartite graph admits a unicyclic spanning subgraph that contains a unique odd cycle; we refer to such a subgraph as an odd unicyclic graph.
Let $H$ be an odd unicyclic graph with $n$ vertices, where $C$ is the unique cycle of $H$. Let $k=|V(C)|$ denote the number of vertices on $C$. We give a labeling of $V(H)$ that satisfies the FM condition.
Step 1: Define the distance from a vertex to cycle $C$. For any vertex $v\in V(H)$, the distance from $v$ to $C$, denoted by $d(v,C)$, is defined recursively as follows:
(1) If $v\in V(C)$, then $d(v,C)=0$ (vertices on the cycle are at distance 0 from $C$);
(2) If $v\notin V(C)$, then $d(v,C)=1+\min\left\{d(u,C)\mid u\in N(v)\right\}$, where $N(v)$ is the set of neighbors of $v$ in $H$.
Step 2: Label vertices in descending order of their distance to $C$. Assign the $k$ largest labels in $[n]$, namely $n,n-1,\dots,n-k+1$, to the vertices of $C$ following a consecutive cyclic order. Then for $t=1,2,\dots$, assign the largest remaining unassigned labels arbitrarily to the vertices in the $t$-th distance layer, proceeding sequentially to the $(t+1)$-th distance layer only after the $t$-th layer is fully labeled. Repeat this process until all vertices of $H$ are labeled.
Lemma 4.12.
Let $H$ be an odd unicyclic graph, where $C$ is the unique cycle of $H$. Label $V(H)$ as the above way. Then this labelling satisfies the FM condition: for any $i\in[n-1]$, the induced subgraph of $H$ on $\{i,i+1,\dots,n\}$ is connected.
Proof.
Fix $i\in[n-1]$. By induction, we may assume that the induced subgraph of $H$ on $[i+1,n]$ is connected. To prove that the induced subgraph of $H$ on $[i,n]$ is connected, it suffices to prove that there exists $i<j\leq n$ such that $i$ is adjacent to $j$. We consider the following two cases.
If $i\geq n-k+1$, then $i\in V(C)$. By the consecutive cyclic labelling rule for $C$ and the fact that $i\leq n-1$, vertex $i$ is adjacent to vertex $i+1$. Thus, we are done.
If i<nk+1i<n-k+1, then $i\notin V(C)$ and so $d(i,C)\geq 1$. By the definition of the distance function, there exists a neighbor $j$ of vertex $i$ such that d(j,C)=d(i,C)1d(j,C)=d(i,C)-1. By the labelling rule, it follows that $j>i$. This completes the proof.
Theorem 4.13.
Let $G$ be a connected non-bipartite graph with $n$ vertices. Then
$\mathrm{dstab}(I_{c}(G))\leq n-2.$
Proof.
Let $H$ be a spanning subgraph of $G$ that is odd unicyclic, and let $C$ be the unique cycle of $H$. Set $k:=|V(C)|$. By the preceding discussion, there exists a labeling of $V(H)$ by $[n]$ that satisfies the FM condition, where the vertices of $C$ are labeled by $n,n-1,\ldots,n-k+1$ in a consecutive cyclic order. Since $H$ is a spanning subgraph of $G$, this labeling also serves as a labeling of $G$ satisfying the FM condition, so Lemma 4.3 is applicable here.
Recall that an odd unicyclic graph has exactly as many edges as vertices (i.e., $|E(H)|=|V(H)|=n$), as it is a tree plus one additional edge forming a cycle. Let $e_{1},e_{2},\ldots,e_{n}$ denote all edges of $H$. Without loss of generality, we assume $e_{n}=\{n-1,n\}$ and $e_{n-1}=\{n-k+1,n\}$. Define
$\beta:=\frac{\alpha^{n-2}}{e_{1}e_{2}\cdots e_{n-3}}\in I_{c}(G)^{n-2}.$
We claim that $\mathrm{set}(\beta)=[n-1]$. To verify this claim, we proceed by cases for $p\in[n-1]$:
Case 1: p=n1p=n-1
There exists some index $1\leq i\leq n-3$ (without loss of generality, take $i=1$) such that $e_{1}=\{n-2,n-1\}$. If $k=3$, then n2=nk+1n-2=n-k+1, so n2n-2 is adjacent to $n$. For $k>3$, n2n-2 is even-connected to $n$ with respect to the product $e_{2}\cdots e_{n-3}$ via the walk: $(n-2)-(n-3)-(n-4)-\cdots-1-n$. By Lemma 4.3, it follows that $p=n-1\in\mathrm{set}(\beta)$.
Case 2: $n-k+1\leq p<n-1$
Since $p\in V(C)$ (by the labeling rule for $C$), we may assume $e_{1}=\{p,p+1\}$. Note that $p+1$ is adjacent to $p+2$, so by Lemma 4.3, $p\in\mathrm{set}(\beta)$.
Case 3: $1\leq p<n-k+1$
Here, $p\notin V(C)$, so by the FM condition for the labeling of $H$, there exists a vertex $q>p$ such that $\{p,q\}\in E(H)$. Suppose first that $q=n$. Set $e_{1}=\{p,q\}=\{p,n\}$. Under this edge designation, since $p\in V(C)$, $e_{1}\in E(H)\setminus E(C)$. Hence, $n$ is even-connected to itself with respect to the product $e_{2}\cdots e_{n-3}$. If $q<n$, then by the FM condition (applied to vertex $q$), there exists a vertex $r>q$ such that $\{q,r\}\in E(H)$. In either case, Lemma 4.3 implies $p\in\mathrm{set}(\beta)$.
This completes the proof of the claim that $\mathrm{set}(\beta)=[n-1]$.
Since $\mathrm{set}(\beta)=[n-1]$, we have $\mathrm{pd}(I_{c}(G)^{n-2})=n-1$. Recall that the projective dimension $\mathrm{pd}(I_{c}(G)^{s})$ is a non-decreasing function of $s\geq 1$, and its maximal possible value is n1n-1. Thus, $\mathrm{pd}(I_{c}(G)^{s})=n-1$ for all $s\geq n-2$, which implies $\mathrm{dstab}(I_{c}(G))\leq n-2$. This completes the proof of the theorem.
5. Homological shift algebra
In this section, we first show that Conjecture 1.1 holds for the complementary edge ideal $I$ of any connected graph when $i=1$. We then establish explicit closed-form expressions for $\mathrm{HS}_{i}(I^{s})$ corresponding to the complementary edge ideals $I$ of trees and cycles. These expressions, in turn, provide a direct proof that Conjecture 1.1 holds for these classes of ideals. We use $\alpha$ to denote ${\bf x}_{[n]}$ in this section.